# Can GenAI Move from Individual Use to Collaborative Work? Experiences, Challenges, and Opportunities of Integrating GenAI into Collaborative Newsroom Routines

### v2

@@ 2.3. GenAI for Teamwork and Organizational Collaboration
Before the rise of GenAI, a long line of research in HCI, CSCW and organizational studies examined how collaborative technologies, from email, shared document platforms, and groupware to chat tools, reshape workplace coordination and communication (Rama and Bishop, 2006; Rodden, 1991; Kong et al., 2024; Carasik and Grantham, 1988; Wang et al., 2022). These tools were primarily designed to improve information flow, enable distributed collaboration, and support decision-making across teams (Cherbonnier et al., 2025; Wang et al., 2022; Handel and Herbsleb, 2002). Early AI systems in organizations, such as expert systems or decision-support tools, were typically domain-specific and structured around narrow tasks (Billi and Labraña, 2025; El-Najdawi and Stylianou, 1993; Abdolmohammadi, 1987; Phillips-Wren, 2012). While useful, these systems rarely participated in open-ended dialogue or generative tasks. In contrast, GenAI introduces a new paradigm: tools that can autonomously generate, reframe, and negotiate meanings within collaborative work. This shift not only expands the scope of automation but also raises new questions about agency and the evolving role of human-AI collaboration (Leong et al., 2024; Shao et al., 2025; (2024), in partnership with Oracle).
@@
According to Gartner’s 2024 survey (Gartner, 2024), GenAI has rapidly become the most widely adopted AI category in organizations. GenAI has opened up new opportunities for teamwork and organizational collaborations. At a basic level, GenAI is already widely used to support routine collaborative functions such as summarizing meeting notes, drafting shared documents, generating ideas, or providing quick answers and even feedback during group discussions. These functions help teams save time, reduce repetitive work, and establish a shared informational baseline (Palani and Ramos, 2024; Han et al., 2024; Takaffoli et al., 2024; Stray et al., 2025; Li et al., 2024a; Liang et al., 2024). More recently, however, research has pointed to emerging developments that extend GenAI’s role from a support tool to an active collaborator. For example, Leong (Leong et al., 2024) introduced the concept of Dittos, AI agents designed as personalized stand-ins that visually and aurally resemble humans and embody their knowledge. These clone agents were found to hold significant potential in team meetings, where they could evoke trust and a sense of presence while providing relevant information to support collective decision-making (Johnson et al., 2025).
Yet enthusiasm is tempered by significant barriers. Organizations struggle to demonstrate clear returns on AI, face shortages of skilled personnel, and encounter persistent data challenges such as access, privacy, and bias (Taherizadeh and Beaudry, 2023; Kleiman, 2025; Gartner, 2024). Organizational misalignment further complicates adoption: siloed initiatives and weak coordination hinder scaling, while knowledge management systems remain fragmented and difficult to build (Zhang et al., 2025b; Dodds et al., 2024). Concerns about trust, transparency, and legitimacy also create internal resistance among employees (Shao et al., 2025; Xiao et al., 2025b; Dodds et al., 2025).
@@ 3. Background: GenAI in the Chinese News Industry
In China, the news industry represents a distinctive setting due to its combination of rapid technological experimentation and strong regulation. The Chinese government has encouraged newsrooms to experiment with GenAI, including initiatives such as developing in-house GenAI systems through continuous recruitment of AI specialists and collaborations with technology companies and university labs (Xiao et al., 2025c; China Press, Publication, Radio, Film and Television Journal, 2024), reflecting a broader ambition to position China at the forefront of AI-driven media transformation. Meanwhile, Chinese newsrooms face significant accountability pressures: errors or misuses of GenAI can result not only in reputational damage among the public but also in institutional repercussions from government authorities (Xiao et al., 2025c). This provides an excellent case for examining how GenAI may potentially be integrated into collaborative newsroom routines.
@@
The Chinese news industry has made notable progress in the design of the GenAI system for news work. [^1] Yet, it is worth noting that the development of these newsroom-specific GenAI systems has not been particularly transparent. For example, most Chinese newsrooms have not disclosed whether their “proprietary” models are entirely self-developed or fine-tuned from existing commercial or open-source foundations. All descriptions summarized here in the background section are directly from the state news reports about Chinese GenAI development in newsrooms, which limits independent verification of the underlying technical processes in this paper. At the organizational level, major media organizations have launched pilot projects that incorporate GenAI into routine news tasks. In July 2023, Xinhua News Agency (China Press, Publication, Radio, Film and Television Journal, 2024) released its proprietary large language model. The model was soon applied in coverage of the Chengdu Universiade, where Xinhua combined GenAI-based image and text generation to create and publish AIGC news videos (China Press, Publication, Radio, Film and Television Journal, 2024). Around the same time, China Media Group (CMG) introduced the “CMG Media GPT,” the first large-scale AI model in China dedicated specifically to audiovisual content production (China Press, Publication, Radio, Film and Television Journal, 2024; Lang, 2024). Also in February 2024, Jiefang Daily unveiled its own GenAI-powered application that could automatically reference the newspaper’s editorial database, and then produces articles in its distinctive newspaper style (China Press, Publication, Radio, Film and Television Journal, 2024). Even regional newspapers also accelerated their investment in domain-specific GenAI models. Zhejiang Radio and Television Group launched an AIGC-based platform to support professional news production and used it to publish news documentaries (China Press, Publication, Radio, Film and Television Journal, 2024). Jiangxi Daily has also developed a GenAI platform and could generate digital AI journalists for reporting (Feng, 2024).
@@
These developments have drawn considerable attention in China’s media policy discourse. As noted by the China Press, Publication, Radio, Film and Television Journal (China Press, Publication, Radio, Film and Television Journal, 2024), which reflects the government’s stance and is directly overseen by the State Administration, such GenAI experiments highlight the need for journalists to develop new paradigms of human–AI collaboration in their daily newswork. Meanwhile, regulatory oversight in China has intensified. From January 2020, producing deceptive or misleading content with AI was made a criminal offense, aimed at deterring unethical applications of the technology (Xiao et al., 2025b). In 2023 the authorities released new rules governing the use of AI in news, which combined stricter accountability with incentives for news organizations to expand the use of GenAI in content production (of China, 2023). Starting from September 1, 2025, the Chinese government requires that all AI-generated content be clearly labeled. Any organization or individual found to maliciously delete, alter, forge, or conceal such labels will be considered in violation of the regulation and subject to penalties (CCTV News, 2025). It is worth noting that the development and application of GenAI in the news industry reflects broader state ambitions to position AI as a driver of technological leadership and cultural influence (Ding, 2018; Zeng, 2020).
@@ 4. Method
To investigate how GenAI is being integrated into collaborative newsroom routines, we conducted 27 in-depth interviews with senior newsroom managers, editors, and front-line journalists in China. All participants were affiliated with news organizations that had initiated early-stage projects involving the development or deployment of GenAI tools. Each journalist had direct experience with GenAI in their collaborative newswork. In the following, we outline our participant recruitment strategy, the design and conduct of the interviews, and our approach to data analysis.
@@ 4.1. Participants
Table 1. Participant Demographics.
ID Gender Age Journalism Experience Position GenAI Experience
@@
P1 Female 48 22 years Editor/Senior Manager 1, 2
P2 Female 42 19 years Editor/Senior Manager 1
@@
P3 Male 45 23 years Editor/Senior Manager 1
P4 Female 38 13 years Editor 1
P5 Female 34 11 years Editor 1, 2
P6 Female 33 10 years Editor 2
P7 Male 37 10 years Editor 1, 2
P8 Male 35 12 years Editor 2
P9 Male 29 3 years Editor 2
P10 Male 28 3 years Front-line Journalist 2
@@
P11 Female 26 2 years Front-line Journalist 2
@@
P12 Female 24 2 years Front-line Journalist 2
@@
P13 Male 32 8 years Editor/Front-line Journalist 2
@@
P14 Female 30 6 years Editor/Front-line Journalist 2
P15 Female 28 3 years Front-line Journalist 1, 2
P16 Female 26 2 years Front-line Journalist 2
@@
P17 Male 25 1 year Front-line Journalist 2
@@
P18 Male 28 3 years Editor/Front-line Journalist 1, 2
@@
P19 Male 25 1 year Front-line Journalist 2
P20 Male 25 1 year Front-line Journalist 2
P21 Male 25 1 year Front-line Journalist 2
P22 Male 25 2 years Front-line Journalist 2
P23 Male 26 1 year Front-line Journalist 2
@@
P24 Female 26 2 years Front-line Journalist 2
P25 Female 26 1 year Front-line Journalist 2
P26 Female 25 1 year Front-line Journalist 2
P27 Male 25 1 year Front-line Journalist 2
@@
Note: GenAI experience codes are defined as follows: 1 = Participation in organizational GenAI strategy or development planning; 2 = Use of GenAI tools in daily newswork routines.
@@
To enrich our data with diverse perspectives and levels of organizational embedding, we sampled across different seniority levels, including editors, senior managers, and front-line journalists. Given our research goal of understanding how GenAI could be integrated into collaborative newsroom routines, we purposely recruited participants who had already engaged with GenAI in their professional roles. Our aim was to examine how journalists with firsthand experience actively negotiate GenAI’s role in teamwork and editorial coordination. While this strategy may bias the sample toward more AI-engaged participants, such a focus is appropriate for capturing the situated practices, tensions, and aspirations surrounding GenAI’s integration into collaborative newswork. Rather than studying those unfamiliar with GenAI and never thinking about future collaborative newswork around GenAI, we sought insight from those already reflecting on its collaborative implications.
17 participants were recruited through the authors’ personal networks, drawing on prior experience in journalism in China, while the remaining 10 were recruited through snowball sampling based on participants’ professional contacts. This approach allowed us to reach a relatively new and specialized population by leveraging trusted professional networks. These participants were not only early adopters of GenAI, but also embedded in organizational environments at the forefront of GenAI innovation. This recruitment strategy is especially appropriate for studying emerging practices that are not yet widely distributed or publicly documented, where recruitment through general channels would be difficult or ineffective (Dicce and Ewers, 2021).
@@
Participants’ journalism experience ranged from one to over twenty years, and they represented a spectrum of positions including front-line journalists (n=16), editors (n=8), and senior managers in newsrooms (n=3). Some participants simultaneously held multiple roles within their organizations, as shown in Table 1. Front-line journalists offered grounded insights into how GenAI tools affect day-to-day reporting in collaborations. Mid-level editors, positioned between strategic directives and ground-level implementation, shared perspectives on coordinating GenAI integration into team workflows and balancing technological adoption with editorial standards. Senior managers contributed a broader organizational view, describing strategic rationales behind GenAI adoption, concerns about institutional credibility, and long-term visions for human–AI collaboration in journalism. Some editors also used GenAI to support their editorial tasks, and some front-line journalists occasionally participated in organizational discussions around GenAI strategy.
To protect participant confidentiality, all individuals were anonymized and assigned pseudonyms in the form of participant IDs ranging from P1 to P27. All participants provided informed consent prior to the interviews, and the study received ethics approval from our institutional review board.
@@ 4.2. Study Design
To explore how journalists experience, negotiate, and reflect on the integration of GenAI into their teamwork routines, we conducted 27 semi-structured interviews. Each interview lasted approximately 90–120 minutes and was conducted remotely. The study was approved by our university’s IRB, and all participants provided informed consent.
Each interview was structured into three phases. First, participants were asked to recall and describe concrete experiences with GenAI in their newsroom work, with particular attention to how these experiences unfolded in collaborative contexts. Participants reflected on how GenAI fit (or failed to fit) into broader editorial routines and teamwork. They were also asked to comment on their organization’s broader AI strategy, if any, and how early-stage GenAI practices were coordinated across the newsroom.
Second, the interview shifted to expectations. We asked participants to imagine what ideal GenAI integration would look like in collaborative newswork. Example questions included: “In your ideal scenario, how would GenAI assist not only you but also your team?” and “What would make a GenAI tool feel reliable in teamwork?”
Finally, participants were invited to reflect critically on the gap between their collaborative ideals and current realities. We asked them to identify what prevents GenAI from being integrated into team workflows, what forms of resistance or uncertainty arise when colleagues work with GenAI, and what organizational or technical changes might better support collective integration. Example questions included: “Why do you think GenAI has not yet become a regular part of your team’s workflow?” and “What organizational or technological changes would make GenAI more compatible with your newsroom’s collaborative practices?”

### v3

@@ 2.3. GenAI for Teamwork and Organizational Collaboration
Before the rise of GenAI, a long line of research in HCI, CSCW and organizational studies examined how collaborative technologies, from email, shared document platforms, and groupware to chat tools, reshape workplace coordination and communication (Rama and Bishop, 2006; Rodden, 1991; Kong et al., 2024; Carasik and Grantham, 1988; Wang et al., 2022). These tools were primarily designed to improve information flow, enable distributed collaboration, and support decision-making across teams (Cherbonnier et al., 2025; Wang et al., 2022; Handel and Herbsleb, 2002). Early AI systems in organizations, such as expert systems or decision-support tools, were typically domain-specific and structured around narrow tasks (Billi and Labraña, 2025; El-Najdawi and Stylianou, 1993; Abdolmohammadi, 1987; Phillips-Wren, 2012). While useful, these systems rarely participated in open-ended dialogue or generative tasks. In contrast, GenAI introduces a new paradigm: tools that can autonomously generate, reframe, and negotiate meanings. This shift not only expands the scope of automation but also raises new questions about agency and the evolving role of human-AI collaboration (Leong et al., 2024; Shao et al., 2025; partnership with Oracle), 2024).
@@
According to Gartner’s 2024 survey (Gartner, 2024), GenAI has rapidly become the most widely adopted AI category in organizations. GenAI has opened up new opportunities for teamwork and organizational collaborations. At a basic level, GenAI is already widely used to support collaborative functions such as summarizing meeting notes, drafting shared documents, generating ideas, or providing quick answers and even feedback during group discussions. These functions help teams save time, reduce repetitive work, and establish a shared informational baseline (Palani and Ramos, 2024; Han et al., 2024; Takaffoli et al., 2024; Stray et al., 2025; Li et al., 2024; Liang et al., 2024). More recently, however, research has pointed to emerging developments that extend GenAI’s role from a support tool to an active collaborator. For example, Leong (Leong et al., 2024) introduced the concept of Dittos, AI agents designed as personalized stand-ins that visually and aurally resemble humans and embody their knowledge. These clone agents were found to hold significant potential in team meetings, where they could evoke trust and a sense of presence while providing relevant information to support collective decision-making (Johnson et al., 2025; Hu et al., 2025).
2.4. Emerging Technologies, Organizational Adoption, and the Limits of Individual Use
A long tradition in CSCW and HCI shows that emerging technologies do not remain individual for long. Although new tools often enter workplaces through personal experimentation, these practices invariably spill over into collaborative norms (Grudin, 1988; Beane, 2019; Cha and Wong, 2025). Technologies such as email, instant messaging, shared document platforms began as “personal tools,” yet their informal uptake had organizational consequences well before formal policies were introduced (Wallace, 2004; Anders, 2016; Bhat et al., 2021).
This scholarship demonstrates that individual use is never merely individual in collaborative environments. When collaborators rely on different tools, use them unevenly, or conceal their usage altogether, articulation work increases and invisible repair labor emerges (Jarrahi and Sawyer, 2015; Grudin, 1988). Even non-use plays an active role: strategic refusal, avoidance, and conditional use express professional values and risk perceptions, which then shape collective workflows and expectations (Selwyn et al., 2005; Garg and Sengupta, 2019; Cha and Wong, 2025; Baumer et al., 2013). As prior studies show, these heterogeneous stances toward a technology reshape teamwork long before an organization formally acknowledges or regulates the tool (Baba, 1999; Li et al., 2024, 2025a). Research on organizational adoption and governance of emerging technologies further demonstrates that when bottom-up use becomes widespread, organizations could grapple with new questions around visibility, accountability, workflow coordination, and legitimacy (Khovanskaya et al., 2020; Banks et al., 2016; Karunakaran, 2024).
Against this backdrop, GenAI could not be treated as an exception. Like earlier waves of workplace technologies, GenAI initially appears as an individualized, private, or experimental practice (Wu, 2024). Yet because journalism is a deeply interdependent domain, where editorial quality, fact-checking, and organizational reputation rely on coordinated judgment, individual GenAI use inevitably produces collective effects. Hidden or uneven GenAI use alters how stories are drafted, how decisions are made, and how outputs are evaluated. It also amplifies well-documented challenges around visibility, provenance, and accountability (Wu, 2024).
Therefore, this study follows directly from decades of HCI and CSCW research demonstrating that once technologies become entangled with collaborative work, questions about governance, transparency, and shared practices necessarily arise. Integrating GenAI into this tradition allows us to see that the core issue is not whether individual journalists use GenAI, but how these individual practices reshape organizational norms and create new forms of interdependence that require collective management.
@@ 3. Background: GenAI in the Chinese News Industry
In China, the news industry represents a distinctive setting due to its combination of rapid technological experimentation and strong regulation. The Chinese government has encouraged newsrooms to experiment with GenAI, including initiatives such as developing in-house GenAI systems through continuous recruitment of AI specialists and collaborations with technology companies and university labs (Xiao et al., 2025d; China Press, Publication, Radio, Film and Television Journal, 2024), reflecting a broader ambition to position China at the forefront of AI-driven media transformation. At the same time, even as organizations pursue large-scale GenAI system development (Xiao et al., 2025d; China Press, Publication, Radio, Film and Television Journal, 2024), individual journalists have begun adopting GenAI tools in their daily practice (Liu, 2025), creating a parallel bottom-up trajectory of experimentation that coexists with top-down infrastructural ambitions.
@@
The Chinese news industry has made notable progress in the design of the GenAI system for news work. [^1] Yet, it is worth noting that the development of these newsroom-specific GenAI systems has not been particularly transparent. For example, most Chinese newsrooms have not disclosed whether their “proprietary” models are entirely self-developed or fine-tuned from existing commercial or open-source foundations. All descriptions summarized here in the background section are directly from the state news reports about Chinese GenAI development in newsrooms, which limits independent verification of the underlying technical processes in this paper. In July 2023, Xinhua News Agency (China Press, Publication, Radio, Film and Television Journal, 2024) released its proprietary large language model. The model was soon applied in coverage of the Chengdu Universiade, where Xinhua combined GenAI-based image and text generation to create and publish AIGC news videos (China Press, Publication, Radio, Film and Television Journal, 2024). Around the same time, China Media Group (CMG) introduced the “CMG Media GPT,” the first large-scale AI model in China dedicated specifically to audiovisual content production (China Press, Publication, Radio, Film and Television Journal, 2024; Lang, 2024). Even regional newspapers also accelerated their investment in domain-specific GenAI models. Zhejiang Radio and Television Group launched an AIGC-based platform to support professional news production and used it to publish news documentaries (China Press, Publication, Radio, Film and Television Journal, 2024).
@@
These developments have drawn considerable attention in China’s media policy discourse. As noted by the China Press, Publication, Radio, Film and Television Journal (China Press, Publication, Radio, Film and Television Journal, 2024), which reflects the government’s stance and is directly overseen by the State Administration, such GenAI experiments highlight the need for journalists to develop new paradigms of human–AI collaboration in their daily newswork. Meanwhile, regulatory oversight in China has intensified. In 2023 the authorities released new rules governing the use of AI in news, which combined stricter accountability with incentives for news organizations to expand the use of GenAI in content production (of China, 2023). Starting from September 1, 2025, the Chinese government requires that all AI-generated content be clearly labeled. Any organization or individual found to maliciously delete, alter, forge, or conceal such labels will be considered in violation of the regulation and subject to penalties (CCTV News, 2025).
@@ 4. Method
Our study investigates how individually adopted GenAI practices begin to intersect with team coordination and collaborative newswork, even when organization-level coordination remains in its early stages. We conducted 27 in-depth interviews with senior newsroom managers, editors, and front-line journalists in China. All participants were affiliated with news organizations that had initiated early-stage projects involving the development or deployment of GenAI tools. Each journalist had direct experience with GenAI in their collaborative newswork. In the following, we outline our participant recruitment strategy, the design and conduct of the interviews, and our approach to data analysis.
@@ 4.1. Participants
Table 1. Participant Demographics
ID Gender Age Journalism Experience Position GenAI Experience GenAI Use Frequency
P1 Female 48 22 years Editor/Senior Manager 1, 2 Several times per day
@@
P2 Female 42 19 years Editor/Senior Manager 1 Several times per week
@@
P3 Male 45 23 years Editor/Senior Manager 1 Several times per week
P4 Female 38 13 years Editor 1 Several times per week
P5 Female 34 11 years Editor 1, 2 Several times per day
P6 Female 33 10 years Editor 2 Several times per week
P7 Male 37 10 years Editor 1, 2 Several times per day
P8 Male 35 12 years Editor 2 Several times per week
P9 Male 29 3 years Editor 2 Several times per week
P10 Male 28 3 years Front-line Journalist 2 Several times per day
@@
P11 Female 26 2 years Front-line Journalist 2 Several times per week
@@
P12 Female 24 2 years Front-line Journalist 2 Several times per week
@@
P13 Male 32 8 years Editor/Front-line Journalist 2 Several times per day
@@
P14 Female 30 6 years Editor/Front-line Journalist 2 Several times per day
P15 Female 28 3 years Front-line Journalist 1, 2 Several times per day
P16 Female 26 2 years Front-line Journalist 2 Several times per day
@@
P17 Male 25 1 year Front-line Journalist 2 Several times per week
@@
P18 Male 28 3 years Editor/Front-line Journalist 1, 2 Several times per day
P19 Male 25 1 year Front-line Journalist 2 Several times per day
P20 Male 25 1 year Front-line Journalist 2 Several times per day
@@
P21 Male 25 1 year Front-line Journalist 2 Several times per week
P22 Male 25 2 years Front-line Journalist 2 Several times per day
P23 Male 26 1 year Front-line Journalist 2 Several times per day
@@
P24 Female 26 2 years Front-line Journalist 2 Several times per week
P25 Female 26 1 year Front-line Journalist 2 Several times per day
P26 Female 25 1 year Front-line Journalist 2 Several times per day
P27 Male 25 1 year Front-line Journalist 2 Several times per day
@@
Note: GenAI experience codes: 1 = Participation in organizational GenAI strategy or development planning; 2 = Use of GenAI tools in daily newswork.
@@
To enrich our data with diverse perspectives and levels of organizational embedding, we sampled across different seniority levels, including editors, senior managers, and front-line journalists. Given our research goal of understanding how GenAI could be integrated into collaborative newswork, we purposely recruited participants who had already engaged with GenAI in their professional roles. Rather than studying those unfamiliar with GenAI and never thinking about future collaborative newswork around GenAI, we sought insight from those already reflecting on its collaborative implications. 17 participants were recruited through the authors’ personal networks, drawing on prior experience in journalism in China, while the remaining 10 were recruited through snowball sampling based on participants’ professional contacts. These participants were not only early adopters of GenAI, but also embedded in organizational environments at the forefront of GenAI innovation.
@@
Participants’ journalism experience ranged from one to over twenty years, and they represented a spectrum of positions including front-line journalists (n=16), editors (n=8), and senior managers in newsrooms (n=3). Some participants simultaneously held multiple roles within their organizations, as shown in Table 1. Front-line journalists offered grounded insights into how GenAI tools affect day-to-day reporting in collaborations. Mid-level editors, positioned between strategic directives and ground-level implementation, shared perspectives on coordinating GenAI into team workflows and balancing technological adoption with editorial standards. Senior managers contributed a broader organizational view, describing strategic rationales behind GenAI coordination, concerns about institutional credibility, and long-term visions for human–AI collaboration in journalism. Some editors also used GenAI to support their editorial tasks, and some front-line journalists occasionally participated in organizational discussions around GenAI strategy.
All participants were drawn from regional news organizations in China, where digital news production has become a central component of daily editorial work. Their newsrooms operate in a dual landscape: on one hand, journalists engage in bottom-up, individualized experimentation with GenAI; on the other hand, organizations are investing heavily in top-down development of proprietary AI tools. Yet despite these parallel trajectories, no formal mechanisms currently exist to coordinate individual GenAI use within collaborative team workflows.
Across these newsrooms, participants regularly engaged in digitally mediated reporting, editing, multimedia production, and platform content distribution. Their primary collaborators included editors responsible for narrative framing and quality assurance, fellow front-line journalists who jointly covered events or exchanged source materials, and, in some organizations, in-house AI engineers and product teams who developed or maintained emerging GenAI tools. Although specific collaboration structures varied by newsroom size and workflow arrangements, most participants worked within small reporting teams or editor–journalist pairs, coordinated through shared content management systems, group chats, and daily editorial meetings. In several cases, participants also had overlapping working relationships, for example, junior journalists who collaborated with multiple editors, or editors who simultaneously supervised and co-produced content with different reporting teams.
Participation in the study was entirely voluntary, and no financial or material compensation was provided, as participants chose to participate based on their genuine interest in GenAI’s implications for journalism rather than external rewards. To protect participant confidentiality, all individuals were anonymized and assigned pseudonyms in the form of participant IDs ranging from P1 to P27.
@@ 4.2. Study Design
To explore how journalists experience, negotiate, and reflect on the coordination of GenAI into their teamwork, we conducted 27 semi-structured interviews. Each interview lasted approximately 90–120 minutes and was conducted remotely. A complete interview protocol, including phase-by-phase guiding questions, is provided in the Appendix A for reference.
Each interview was structured into three phases. First, participants were asked to recall and describe concrete experiences with GenAI in their newsroom work, with particular attention to how these experiences unfolded in collaborative contexts. Second, the interview shifted to expectations. We asked participants to imagine what ideal GenAI coordination would look like in collaborative newswork. Example questions included: “In your ideal scenario, how would GenAI assist not only you but also your team?” and “What would make a GenAI tool feel reliable in teamwork?” Finally, participants were invited to reflect critically on the gap between their collaborative ideals and current realities. We asked them to identify what prevents GenAI from being coordinated into team workflows, and what organizational or technical changes might better support GenAI in collaborative newswork.